### v2

# TopoBDA: Towards Bezier Deformable Attention for Road Topology Understanding

Abstract
Understanding road topology is crucial for autonomous driving. This paper introduces TopoBDA (Topology with Bezier Deformable Attention), a novel approach that enhances road topology comprehension by leveraging Bezier Deformable Attention (BDA). TopoBDA processes multi-camera 360-degree imagery to generate Bird’s Eye View (BEV) features, which are refined through a transformer decoder employing BDA. BDA utilizes Bezier control points to drive the deformable attention mechanism, improving the detection and representation of elongated and thin polyline structures, such as lane centerlines. Additionally, TopoBDA integrates two auxiliary components: an instance mask formulation loss and a one-to-many set prediction loss strategy, to further refine centerline detection and enhance road topology understanding. Experimental evaluations on the OpenLane-V2 dataset demonstrate that TopoBDA outperforms existing methods, achieving state-of-the-art results in centerline detection and topology reasoning. TopoBDA also achieves the best results on the OpenLane-V1 dataset in 3D lane detection. Further experiments on integrating multi-modal data—such as LiDAR, radar, and SDMap—show that multimodal inputs can further enhance performance in road topology understanding.
Keywords:
Road Topology Understanding Centerline Detection Autonomous Driving Automated HDMap Generation 3D Lane Detection
1 Introduction
In autonomous driving, scenes primarily consist of two types of entities: dynamic and stationary. Dynamic entities encompass objects capable of movement and interaction, such as vehicles (cars, bicycles, motorcycles, trucks), and pedestrians. Stationary entities, on the other hand, are immobile objects and abstract structures such as lane markings, crosswalks, road signs, traffic lights, barriers, and road surfaces, that regulate traffic and ensure the safe and orderly movement of dynamic entities. Furthermore, abstract stationary entities -such as lanes and centerlines- are defined in relation to other stationary objects or specific rules of the driving environment.
For a fully autonomous driving system, it is essential not only to detect stationary entities but also to understand their interrelationships. The challenge of identifying and understanding these connections between stationary objects is referred to as the road topology problem. For instance, multiple lanes may converge into one or a few, or a single lane may split into several. This complexity increases at intersections, where numerous lanes interact. Additionally, some traffic lights control only specific lanes. Accurately localizing, categorizing, and understanding the relationships between stationary entities are essential for downstream tasks such as planning and control in autonomous driving systems.
An alternative solution to this problem is the utilization of High-Definition Maps (HDMaps), which provide pre-computed maps of stationary entities. However, HDMaps are expensive to produce, cover only limited geographic areas, and cannot reflect recent changes on the road, requiring continuous updates. Furthermore, errors introduced by the Global Navigation Satellite System (GNSS) receiver on the vehicle can lead to localization inaccuracies. These inaccuracies may cause discrepancies between the actual position of the vehicle and the HDMap data, potentially resulting in drifts in the map-based guidance system. In response to these challenges, automatic HDMap construction has gained significant attention in recent years for two primary reasons [26, 39, 34, 3]. First, it can reduce reliance on HDMaps for autonomous driving. Second, it lowers the cost of creating and maintaining HDMaps, thereby minimizing the need for human effort.
Figure 1: Comparison of various cross-attention mechanisms within the decoder architecture for polyline structures.
In the context of polyline structures, such as centerlines and lane dividers, the application of standard cross attention [3, 40], deformable cross attention [27, 66, 34, 46, 28, 39, 1, 82, 9, 42, 70, 76], and masked cross attention [53, 10, 82, 21] is common. Masked attention [8], as illustrated in Figure 1b, requires tuning hyperparameters for polyline width and uses a thresholding mechanism to differentiate between foreground and background. Moreover, to prevent the failure of the attention mechanism, a foreground check is performed for every query, increasing complexity and reducing deployment efficiency. In contrast, deformable attention [83] can be implemented in two primary ways. The first, Single-Point Deformable Attention (SPDA), shown in Figure 1c, limits attention for each polyline instance to a single point, typically the center point of the polyline’s bounding box [39, 66, 66]. However, this method restricts attention to a local scope, making it unsuitable for elongated, thin polyline structures.
The second method for implementing deformable attention, Multi-Point Deformable Attention (MPDA), as shown in Figure 1d, involves distributing attention around every predicted dense lane point [34, 77, 46, 70, 4]. From the MPDA perspective, there are two primary approaches: point query-based methods and instance query-based methods. Point query-based methods are inherently complex because each query represents a single point, and increasing the number of points per instance proportionally increases the complexity [34, 35, 46, 10, 25]. Conversely, instance query-based methods are more efficient [28, 77]. However, a gap in the literature exists, as instance query methods utilizing Bezier control points [66, 27] have not applied MPDA, relying instead on SPDA.
This study focuses on enhancing centerline detection performance within the broader context of road topology understanding. First, the integration of MPDA into Bezier keypoint-dependent transformer decoder structures is introduced. This adaptation is considered crucial because, unlike other dense polyline prediction methods, the Bezier representation [28] [77], leverages its inherently compact nature to improve the computational efficiency of polyline prediction. By incorporating MPDA into the Bezier keypoint-dependent transformer decoder, the method can more effectively handle elongated and thin polyline structures.
To further optimize the balance between performance and computational complexity, Bezier Deformable Attention (BDA) is introduced. As illustrated in Figure 1e, this method generates deformable attention around predicted Bezier points for centerline prediction. BDA significantly outperforms SPDA while maintaining a negligible performance difference. The improvement, as well as the reason for not observing an increase in computational complexity, is attributed to the use of control points across different attention heads, rather than relying on a single point to drive all attention heads. Additionally, SPDA requires learning an additional regression target for the center of the bounding box of the centerline. Furthermore, BDA achieves slightly better performance than the MPDA adaptation to the Bezier concept, with slightly less computational overhead. Unlike MPDA, BDA eliminates the need to convert Bezier keypoints into multiple polyline points within each transformer decoder layer, thereby reducing computational complexity.
In addition, consistent with the findings of the TopoMaskV2 study [21], the instance-mask formulation is employed to enhance the overall performance of road topology. However, unlike the direct approach used in TopoMaskV2, TopoBDA adopts an indirect application to reduce post-processing requirements. First, it has been demonstrated that incorporating the instance-mask formulation as an auxiliary loss significantly benefits the Bezier head. Second, during the Hungarian matching, using a Mask-L1 mix matcher [24], instead of a pure L1 matcher, has proven to be superior. These proposed mechanisms underscore the effectiveness of the indirect instance-mask formulation in improving road topology performance.
Sensor fusion has shown significant benefits in various domains, including 3D lane detection [44, 45] and HDMap element prediction [38, 78, 34, 39, 26]. Despite these advancements, there remains a notable gap in the literature regarding its application to road topology understanding. Previous studies have primarily focused on the performance gains of SDMap [43, 74], without exploring the potential of sensor fusion in this context. Our research is the first to investigate and comprehensively evaluate the effects of sensor fusion, utilizing both lidar and radar data for road topology understanding. Additionally, we analyze the benefits of integrating SDMap with lidar and camera sensors, highlighting the novel combination of lidar and SDMap, which has not been explored in the literature.
Auxiliary one-to-many set prediction loss strategy, adapted from hybrid matching techniques [20], is implemented for HDMap element prediction [35] and shown to improve convergence and performance without increasing the inference complexity. While, also employed in road topology understanding problem [66, 21], its quantitative benefits in this context have not been fully explored. Therefore, this research provides a comprehensive analysis of its impact on the proposed TopoBDA architecture.
With the inclusion of BDA, instance mask formulation, and one-to-many set prediction loss, TopoBDA achieves state-of-the-art results in the camera-only benchmark for both OpenLane-V1 and OpenLane-V2 datasets. Furthermore, when utilizing multi-modal data, TopoBDA attains state-of-the-art results in OpenLane-V2. To summarize, the key innovations and contributions of this work are as follows:
  • Multi-Point Deformable Attention (MPDA): The performance of centerline detection and road topology understanding is enhanced by the novel adaptation of MPDA to Bezier keypoint-dependent transformer decoders.
  • Bezier Deformable Attention (BDA): A novel attention mechanism utilizing Bezier control points is introduced. It has been demonstrated that BDA significantly improves the performance of centerline detection and road topology understanding while incurring negligible computational complexity overhead.
  • Instance Mask Formulation: An instance mask formulation is incorporated as an auxiliary loss alongside the Mask-L1 mix matcher, improving the overall performance in road topology understanding.
  • Multi-Modal Fusion: Lidar and radar data are utilized for the first time specifically for road topology understanding. Additionally, the fusion of lidar with SDMap is analyzed, demonstrating the benefits of integrating multi-modal data. By fusing camera, lidar, and SDMap data, state-of-the-art results are achieved.
  • Auxiliary One-to-Many Set Prediction Loss: The auxiliary one-to-many set prediction loss strategy from the existing literature is adapted for the road topology understanding and experimentally evaluated for the first time.
2 Related Work
This section summarizes the literature on three key aspects of autonomous driving and advanced driver-assistance systems (ADAS): Lane Divider Detection, HDMap Element Prediction, and the Road Topology Problem and Centerline Concept. Lane Divider Detection focuses on accurately identifying lane boundaries to ensure safe and reliable vehicle navigation. HDMap Element Prediction, which also encompasses lane divider detection, involves forecasting the presence and attributes of high-definition map elements. These studies often utilize multi-camera setups that provide 360-degree coverage, which is essential for precise localization and path planning. Additionally, the Road Topology Problem and Centerline Concept, both integral to HDMap Element Prediction, address the challenges of understanding and representing the road network’s structure. Together, these components form the backbone of contemporary research aimed at enhancing the safety and efficiency of autonomous vehicles.
2.1 Lane Divider Detection
Lane divider detection methods can be categorized into two primary subcategories: perspective view methods and 3D lane divider methods.
2.1.1 Perspective View Methods
Perspective view (PV) methods focus on detecting lane dividers from the PV and projecting them onto the ground using a homography matrix under the flat surface assumption. Despite the presence of various lane divider instances, semantic approaches are practical and effective, assuming a constant number of lane divider instances (e.g., 1st left, 2nd left, 1st right, 2nd right lane dividers). SCNN [50] adheres to this methodology and introduces a module that sequentially processes the rows and columns of the feature map. UFLD [54] transitions the formulation from pixel-based to grid-based, proposing row-based and column-based anchor formulations to enhance inference speed. LaneATT [56], inspired by object detection studies, devises an anchor concept specifically for lanes. The emergence of the CurveLanes dataset [68] has led to the adoption of instance-segmentation-based methods for the PV domain. CondLaneNet [36] introduces lane-specific methodologies such as offset prediction and row-wise formulation on top of the instance mask formulation in the PV domain. PolyLaneNet [57] and BezierLaneNet [11] employ polynomial and Bezier curve representations, respectively, to reduce post-processing efforts and improve curve learning.
2.1.2 3D Lane Divider Methods
Recent studies have shifted towards directly predicting the 3D locations of lane dividers with the introduction of 3D lane divider datasets such as [72], OpenLane [5], and Apollo 3D Synthetic Lane [13]. This approach addresses the limitations of PV methods, specifically the flat world assumption due to the absence of depth information. Persformer [5] utilizes a deformable attention-based decoder with Inverse Perspective Mapping (IPM) and formulates a 3D anchor concept. CurveFormer [1] employs a sparse query design with a deformable attention mechanism and predicts polynomial parameters in the BEV domain. BEV-LaneDet [61] uses a keypoint concept with predicted offsets and groups the keypoints of the same lane instance with an embedding concept. PETRV2 [40] extends its sparse query design for lane detection. M2-3DLaneNet explores the benefits of integrating lidar and camera sensors. Another study [7] follows an instance-offset formulation in the BEV domain and aggregates offsets with a voting mechanism. LATR [46] introduces an end-to-end 3D lane detection framework that directly detects 3D lanes from front-view images using lane-aware queries and dynamic 3D ground positional embedding, significantly improving accuracy and efficiency over previous methods. GLane3D [84] utilizes a graph-based approach with keypoints and directed connections, achieving enhanced cross-dataset generalization performance.
2.2 HDMap Element Prediction
The prediction of High-Definition Map (HDMap) elements, such as lane dividers, road dividers, and pedestrian crossings, is essential for autonomous driving. Various methods have been developed to enhance the accuracy and efficiency of HDMap element prediction. HDMapNet [26] transforms Bird’s Eye View (BEV) semantic segmentation into BEV instance segmentation through extensive post-processing, leveraging predicted instance embeddings and directional information. VectorMapNet [39] introduces an end-to-end vectorized HD map learning pipeline that predicts sparse polylines from sensor observations, explicitly modeling spatial relationships between map elements, thus eliminating the need for dense rasterized segmentation and heuristic post-processing. Contrary to the autoregressive structure of VectorMapNet, MapTR [34] directly predicts points on polylines or polygons using a permutation-invariant Hungarian matcher, combining point query and instance query concepts, achieving real-time inference speeds and robust performance in complex driving scenes. InstaGraM [55] redefines polyline detection as a graph problem, where keypoints are vertices and their connections are edges, leveraging graph neural networks to enhance accuracy and robustness. MGMap [38] uses instance masks to generate map element queries and refine features with mask outputs, significantly improving performance over baseline methods. MapVR [78] rasterizes MapTR outputs and applies instance segmentation loss to address keypoint-based method limitations, enhancing performance without extra computational cost during inference. ADMap [16] employs instance interactive attention and vector direction difference loss to reduce point sequence jitter, enhancing map accuracy and stability. BeMapNet [53] models map elements as multiple piecewise curves using Bezier curves, eliminating the need for post-processing and achieving superior performance on existing benchmarks. StreamMapNet [77] explores the temporal aspects of HDMap element prediction, using temporal information as propagated instance queries and warped BEV features in a recurrent manner. MapTracker [4] formulates mapping as a tracking task, maintaining multiple memory latents to ensure consistent reconstructions over time, significantly outperforming existing methods on consistency-aware metrics. Recent advancements include PriorMapNet [60], which enhances online vectorized HD map construction by incorporating priors, and HIMap [82], which integrates point-level and element-level information to improve prediction accuracy. Additionally, the Multi-Session High-Definition Map-Monitoring System [64] employs machine learning algorithms to track and update map elements across multiple sessions, ensuring HD maps remain accurate and up-to-date. The Ultra-fast Semantic Map Perception [71] leverages both camera and LiDAR data to achieve real-time performance, featuring an orthogonal projection subspace for fast semantic segmentation and a Bayesian framework for enhanced global semantic fusion.
2.3 Road Topology Problem and Centerline Concept
Road topology refers to the interrelationships among lanes, as well as their connections to traffic lights and signs. However, using lane dividers for this problem is inefficient, as each lane requires two separate lane dividers. Consequently, the concept of centerlines has emerged as a more efficient and natural representation of lanes.
2.3.1 Centerline Concept
STSU [3] introduced a novel approach for extracting a directed graph of the local road network in bird’s-eye-view (BEV) coordinates from a single onboard camera image, significantly improving traffic scene understanding. CenterLineDet [69] uses a transformer network to detect lane centerlines with vehicle-mounted sensors, effectively handling complex graph topologies such as lane intersections. LaneGAP [33] models lane graphs path-wise, preserving lane continuity and improving the accuracy of lane graph construction. MapTRV2 [35] enhances centerline prediction efficiency by treating lane centerlines as paths and incorporating semantic-aware shape modeling. SMERF [43] integrates Standard Definition (SD) maps by tokenizing map elements using a transformer encoder, and employs these tokens in the cross-attention mechanism of a transformer decoder to improve lane detection and topology prediction. In contrast, TopoSD [74] uses both map tokens and feature maps extracted from SD maps to enrich BEV features. SMART [75] uniquely leverages SD and satellite maps to learn robust map priors, enhancing lane topology reasoning for autonomous driving without relying on consistent sensor configurations. LaneSegNet [28] introduces the concept of lane segments, combining geometry and topology information to provide a comprehensive representation of road structures.
2.3.2 Road Topology
TopoNet [27] proposes a graph neural network architecture that models the relationships between centerlines and traffic elements. It incorporates prior relational knowledge to enhance feature interactions. TopoMLP [66] introduces an advanced pipeline for understanding driving topology by incorporating lane coordinates into the topology framework and using an L1 loss function to refine the interaction points. CGNet [14] focuses on preserving the continuity of centerline graphs and improving topology accuracy through modules like Junction Aware Query Enhancement and Bezier Space Connection. Topo2D [25] integrates 2D lane priors to improve 3D lane detection and topology reasoning. TopoLogic [12] proposes managing lane topology relationships by combining geometric lane distance with similarity-based topology relationships. TopoFormer [47] introduces a lane aggregation layer that leverages geometric distance in driving self-attention, along with a counterfactual intervention layer to improve reasoning by considering alternative scenarios and their causal impacts.
3 Methodology
TopoBDA (Figure 2) starts by extracting Bird’s Eye View (BEV) features from multi-camera 360-degree imagery. A transformer decoder then processes these BEV features within its cross-attention mechanism using a sparse query approach. In this approach, each query corresponds to a centerline instance rather than individual points, which improves computational efficiency. Subsequently, the decoder predicts Bezier control points for each instance query which are then converted into dense polyline points via matrix multiplication. The Bezier representation offers a compact and compressed formulation, reducing computational complexity at the regression heads [11, 53, 27]. Additionally, each centerline query predicts an instance mask for auxiliary loss, though these masks are not utilized during inference.
Figure 2: Overview of the TopoBDA architecture. The TopoBDA architecture is based on the instance query concept. The extracted BEV features from the multiple camera images are fed into the transformer decoder. The decoder outputs Bezier control points for each query, which are then converted into centerline instances via matrix multiplication. Additionally, each centerline query predicts instance masks, but only during training.
In the TopoBDA architecture, the BEV feature extraction (Figure 2) starts with a set of $N$ images, $\{\mathbf{I}_{i}\}_{i=1}^{N}$, each having a dimension of $H_{I}\times W_{I}\times 3$, with $H_{I}$ and $W_{I}$ representing the height and width, respectively. These images are first converted into perspective view features, $\{\mathbf{F}_{PV_{i}}\}_{i=1}^{N}$, using a feature extraction function $f_{PV}$, such that $\mathbf{F}_{PV_{i}}=f_{PV}(\mathbf{I}_{i})$ and $\mathbf{F}_{PV_{i}}\in\mathbb{R}^{H_{PV}\times W_{PV}\times C_{PV}}$. These perspective view features are then aggregated and projected into a single Bird’s Eye View (BEV) feature map, $\mathbf{F}_{BEV}$, using a projection function $f_{BEV}$. This function can be implemented via Lift-Splat-Shoot (LSS) [51, 18, 30], transformers [31, 6, 81, 62], or other projection techniques [29, 67, 15]. This projection is defined as $\mathbf{F}_{BEV}=f_{BEV}(\{\mathbf{F}_{PV_{i}}\}_{i=1}^{N})$, where $\mathbf{F}_{BEV}\in\mathbb{R}^{H_{BEV}\times W_{BEV}\times C_{BEV}}$. This methodology effectively captures and transforms spatial features into the BEV space, facilitating further analysis and processing within the TopoBDA framework.
Next, the extracted BEV features $\mathbf{F}_{BEV}$ are fed into the transformer decoder, where they are used as the value input in the cross-attention mechanism. Each query, corresponding to a single centerline instance, interacts with $\mathbf{F}_{BEV}$, and Bezier control points are predicted at each decoder layer. The attention mechanism in TopoBDA is based on Bezier Deformable Attention (BDA) and it is guided by these predicted Bezier control points. Further details of this mechanism are provided in Section 3.1 and the overall structure of the TopoBDA transformer decoder is provided in Section 3.2.
In the transformer decoder of the TopoBDA, an auxiliary instance mask formulation is also employed (Figure 2). In this formulation, each centerline query predicts not only Bezier control points but also a mask probability map for each instance. This indirect utilization enhances the performance of centerline predictions generated from the Bezier control points. Additionally, replacing the pure L1 matcher with a Mask-L1 mix matcher in the Hungarian matcher process further improves accuracy. The details of the instance mask formulation are provided in Section 3.3. This section also introduces the multi-modal data fusion strategy (Section 3.4) and the auxiliary one-to-many set prediction loss strategy (Section 3.5), both of which are integral to the TopoBDA study.
3.1 Towards Bezier Deformable Attention
This section explains the transition from Single-Point Deformable Attention (SPDA) to Multi-Point Deformable Attention (MPDA) for methodologies based on Bezier keypoint regression (or Bezier control point regression). This transition replaces traditional deformable attention heads with dense polyline points obtained via matrix multiplication on Bezier control points (Figure 3a).
Next, the concept of Bezier Deformable Attention (BDA) is introduced. BDA enhances the deformable attention mechanism by using Bezier control points. In this approach, traditional attention heads are substituted with the control points of a polyline curve, facilitating more flexible and adaptive attention. In each layer, the predicted control points are utilized to guide the Bezier Deformable Attention process (Figure 3b).
3.1.1 Bezier Curve Representation
A Bezier curve is a parametric curve frequently used in computer graphics. It is defined by a set of control points, and the curve is a linear combination of these points weighted by Bernstein polynomials.
A Bezier curve $\mathbf{S}(t)$ of order $N$ is defined by $N+1$ control points $\mathbf{C}=\{\mathbf{c}_{0},\mathbf{c}_{1},\ldots,\mathbf{c}_{N}\}$, as shown in Eq. (1), where $B_{n,N}(t)$ are the Bernstein basis polynomials of degree $N$ which are obtained as in Eq. (2).
$\mathbf{S}(t)=\sum_{n=0}^{N}B_{n,N}(t)\mathbf{c}_{n},\quad 0\leq t\leq 1.$ (1)
$B_{n,N}(t)=\binom{N}{n}t^{n}(1-t)^{N-n}.$ (2)
The control points $\mathbf{c}_{n}$ determine the shape of the Bezier curve, and the Bernstein polynomials $B_{n,N}(t)$ provide a smooth interpolation between these points.
3.1.2 Deformable Attention Mechanism
The deformable attention mechanism enhances traditional attention by adaptively sampling reference points with learned offsets, enabling greater flexibility. Single Point Deformable Attention (SPDA) mechanism is defined as in Eq. (3).
$\text{SPDA}(\mathbf{q},\mathbf{V},\mathbf{p})=\sum_{m=1}^{M}\sum_{k=1}^{K}A_{m,k}\mathbf{W}_{m}\mathbf{V}(\mathbf{p}+\Delta\mathbf{p}_{m,k}),$ (3)
where $\mathbf{q}$ is the query, $\mathbf{V}$ is the value matrix, $A_{m,k}$ are the attention weights, $\mathbf{W}_{m}$ are learnable weight matrices, $\mathbf{p}$ is the reference point, $\Delta\mathbf{p}_{m,k}$ are the offsets, $M$ is the number of attention heads, $K$ is the number of sampling points per attention head. In this mechanism, the query $\mathbf{q}$ attends to the value matrix $\mathbf{V}$ at positions determined by the reference point $\mathbf{p}$ and the learned offsets $\Delta\mathbf{p}_{m,k}$.
3.1.3 Adaptation of Bezier Regression Methods to Multi-Point Deformable Attention (MPDA)
Figure 3: Comparison of Multi-Point Deformable Attention (MPDA) and Bezier Deformable Attention (BDA): MPDA necessitates an additional matrix multiplication block within each transformer decoder. Despite their different input utilizations, the mechanisms of MPDA and BDA are fundamentally the same.
This methodology involves extracting $L+1$ polyline points $\mathbf{P}=\{\mathbf{p}_{0},\mathbf{p}_{1},\ldots,\mathbf{p}_{L}\}$ from the control points $\mathbf{C}=\{\mathbf{c}_{0},\mathbf{c}_{1},\ldots,\mathbf{c}_{N}\}$ through Bezier extraction process. These polyline points are then utilized in the deformable attention mechanism. This approach replaces the traditional multi-head attention mechanism, which operates on a single reference point, with a more flexible and adaptive mechanism. Instead of using multiple heads $M$ on a single point $\mathbf{p}$, this method employs a single head for each of the dense polyline points $\mathbf{p}_{l}$. By utilizing multiple points distributed along the polyline, this technique captures the thin, elongated characteristics of polylines, thereby enhancing the attention mechanism’s ability to model complex patterns and dependencies.
The Multi-Point Deformable Attention (MPDA) mechanism is defined as:
$\text{MPDA}(\mathbf{q},\mathbf{V},\mathbf{P})=\sum_{l=0}^{L}\sum_{k=1}^{K}A_{l,k}\mathbf{W}_{l}\mathbf{V}(\mathbf{p}_{l}+\Delta\mathbf{p}_{l,k}),$ (4)
where $L+1$ is the number of polyline points, $K$ is the number of sampling points for each polyline point, $\mathbf{p}_{l}$ represents the polyline points extracted from the Bezier control points and members of $\mathbf{P}$ such that $\mathbf{P}=\{\mathbf{p}_{0},\mathbf{p}_{1},\ldots,\mathbf{p}_{L}\}$.
The polyline points $\mathbf{p}_{l}$ can be extracted from the control points using the following formula, which is derived from Eq. (1) and Eq. (2):
$\mathbf{p}_{l}=\sum_{n=0}^{N}\binom{N}{n}t_{l}^{n}(1-t_{l})^{N-n}\mathbf{c}_{n},$ (5)
for $l=0,1,\ldots,L$, where $t_{l}$ are uniformly spaced within the interval $[0,1]$. Eq. (5) demonstrates that adapting MPDA to the Bezier keypoint concept necessitates converting Bezier control points into polyline points. This conversion is implemented through matrix multiplication in each decoder layer, as illustrated in Figure 3a. The detailed process of this implementation is provided in Section S.1 of the supplementary materials.
3.1.4 Bezier Deformable Attention
BDA enhances the traditional multi-head attention mechanism by replacing the single-point focus $\mathbf{p}$ with control points $\mathbf{c}_{n}$ of the Bezier curve. Unlike MPDA, which relies on dense polyline points, BDA directly uses these control points as reference points within the deformable attention mechanism. This approach eliminates the need for converting control points to polyline points in each decoder layer (Figure 3), reducing computational complexity. By predicting Bezier control points and focusing attention around them, BDA improves the learning process, leading to more effective and accurate predictions.
The Bezier Deformable Attention (BDA) mechanism is defined as:
$\text{BDA}(\mathbf{q},\mathbf{V},\mathbf{C})=\sum_{n=0}^{N}\sum_{k=1}^{K}A_{n,k}\mathbf{W}_{n}\mathbf{V}(\mathbf{c}_{n}+\Delta\mathbf{p}_{n,k}),$ (6)
where $N+1$ is the number of control points, $K$ is the number of sampling points for each control point, $\mathbf{c}_{n}$ represents the control points and members of $\mathbf{C}$ such that $\mathbf{C}=\{\mathbf{c}_{0},\mathbf{c}_{1},\ldots,\mathbf{c}_{N}\}$. In this formulation, each control point $\mathbf{c}_{n}$ acts as a head in the attention mechanism, allowing the model to attend to different parts of the input sequence based on the shape of the Bezier curve. This approach provides a more flexible and adaptive attention mechanism that can better capture the spatial structure of polylines.
3.2 Bezier Deformable Attention Based Transformer Decoder
Figure 4: This figure visualizes the layers of TopoBDA, each driven by Bezier Deformable Attention (BDA) using control points predicted through iterative refinement. Note that iterative refinement is not applicable to the first layer, which uses direct prediction.
The overview of the general pipeline of each layer in the transformer decoder of TopoBDA is shown in Figure 4. The process involves several steps:
  1. 1.
    Control Points Generation (First Layer): The control points of each centerline instance are obtained in a normalized format using a Multi-Layer Perceptron (MLP) and a sigmoid function:
    $\mathbf{C}_{norm}^{(1)}=\sigma(\text{MLP}_{B}^{(1)}(\mathbf{E}_{query}^{(1)})),$ (7)
    where $\mathbf{C}_{norm}^{(1)}=\{\mathbf{c}_{0},\mathbf{c}_{1},\ldots,\mathbf{c}_{N}\}$, and each $\mathbf{c}_{i}$ consists of normalized (scaled to [0,1]) $x$, $y$, and $z$ coordinates. $\mathbf{E}_{query}^{(1)}$ is the query embedding at the first layer.
  2. 2.
    Bezier Deformable Attention (BDA): The predicted control points are fed into the BDA layer to guide attention. BDA defines the query as the sum of positional embedding and query embedding, while the BEV features serve as the value:
    $\mathbf{Q}_{BDA}^{(l)}=\mathbf{E}_{query}^{(l)}+\mathbf{P}^{(l)},$ (8)
    $\mathbf{A}_{BDA}^{(l)}=\text{BDA}(\mathbf{Q}_{BDA}^{(l)},\mathbf{F}_{BEV},\mathbf{C}_{norm}^{(l)}),$
    where $\mathbf{Q}_{BDA}^{(l)}$ is the combined query embedding and positional embedding at layer $l$, and $\mathbf{A}_{BDA}^{(l)}$ is the output of the BDA layer.
  3. 3.
    Self-Attention: Self-attention follows the output of BDA and utilizes the same positional embedding:
    $\mathbf{A}_{SA}^{(l)}=\text{SelfAttention}(\mathbf{A}_{BDA}^{(l)},\mathbf{P}^{(l)}),$ (9)
    where $\mathbf{A}_{SA}^{(l)}$ is the output of the self-attention layer at layer $l$. The output of the self-attention layer, $\mathbf{A}_{SA}^{(l)}$, serves as the query embedding $\mathbf{E}_{query}^{(l+1)}$ for the next layer.
  4. 4.
    Iterative Process in Subsequent Layers: In subsequent layers, the process repeats, but with MLP layers predicting Bezier control points differences. These differences are summed in the inverse sigmoid domain before being transformed back:
    $\Delta\mathbf{C}^{(l)}=\text{MLP}_{B}^{(l)}(\mathbf{E}_{query}^{(l)}),$ (10)
    $\mathbf{C}_{inv\_sigmoid}^{(l)}=\sigma^{-1}(\mathbf{C}_{norm}^{(l-1)})+\Delta\mathbf{C}^{(l)},$
    $\mathbf{C}_{norm}^{(l)}=\sigma(\mathbf{C}_{inv\_sigmoid}^{(l)}),$
    where $\Delta\mathbf{C}^{(l)}$ represents the predicted Bezier control points differences at layer $l$, $\mathbf{C}_{inv\_sigmoid}^{(l)}$ is the sum in the inverse sigmoid domain, and $\mathbf{C}_{norm}^{(l)}$ are the updated control points.
3.3 The Indirect Benefits of Instance Mask Formulation
The TopoBDA architecture is further enhanced by integrating instance mask formulation in two key areas. First, instance mask loss is used as an auxiliary loss. In each transformer decoder layer, instance masks are predicted for each centerline instance. Second, the Mask-L1 mix matcher is introduced in the Hungarian matcher, combining mask and L1 losses to improve the accuracy of predictions by leveraging both mask and L1 bipartite-matching mechanisms.
3.3.1 Instance Mask Formulation as an Auxiliary Loss
Figure 5: The implementation of instance mask formulation in the TopoBDA decoder. Query embeddings are converted to mask embeddings by MLP layers. The dot product between BEV features and mask embeddings generates a pre-mask map, which is iteratively summed across layers to produce the final pre-mask map
To achieve this, centerlines are converted into masks using a parameter $W$ in the BEV domain. Utilizing the generated ground truth instance masks, each centerline query in the transformer decoder predicts both Bezier control points and a mask probability map for each centerline instance (Figure 5).
The mask probability map generation in the decoder follows several steps:
  1. 1.
    Mask Embedding Generation: The mask embeddings are generated from the query embeddings using an MLP:
    $\mathbf{E}_{mask}^{(l)}=\text{MLP}_{M}^{(l)}(\mathbf{E}_{query}^{(l)}),$ (11)
    where $\mathbf{E}_{mask}^{(l)}$ is the mask embedding at layer $l$ and $\mathbf{E}_{query}^{(l)}$ is the query embedding at layer $l$.
  2. 2.
    Pre-mask Map Generation: A dot product is applied between the BEV features ($\mathbf{F}_{BEV}$) and the mask embedding to generate the pre-mask map:
    $\mathbf{P}_{mask}^{(l)}=\mathbf{F}_{BEV}\cdot\mathbf{E}_{mask}^{(l)},$ (12)
    where $\mathbf{P}_{mask}^{(l)}$ is the pre-mask map at layer $l$.
  3. 3.
    Summation of Pre-mask Maps: In consecutive layers, pre-mask maps are summed to refine the mask embeddings at each layer:
    $\mathbf{P}_{mask}^{(l)}=\mathbf{P}_{mask}^{(l-1)}+\mathbf{P}_{mask}^{(l)},$ (13)
    where $\mathbf{P}_{mask}^{(l)}$ is the pre-mask map at layer $l$, $\mathbf{P}_{mask}^{(l-1)}$ is the pre-mask map from the previous layer and $\mathbf{P}_{mask}^{(0)}$ is initialized to zero.
  4. 4.
    Mask Probability Map: The sigmoid function is used to obtain the mask probability maps from the pre-mask maps:
    $\mathbf{M}_{prob}^{(l)}=\sigma(\mathbf{P}_{mask}^{(l)}),$ (14)
    where $\mathbf{M}_{prob}^{(l)}$ is the mask probability map at layer $l$ and $\sigma$ is the sigmoid function.
3.3.2 Mask-L1 Mix Matcher
The instance mask concept is employed not only in the main loss but also during the bipartite matching (Hungarian matcher). The loss function of the bipartite matching is defined as follows:
$\mathcal{L}_{\text{l}}=\lambda_{\text{reg}}\mathcal{L}_{\text{reg}}+\lambda_{\text{mask}}\mathcal{L}_{\text{mask}}+\lambda_{\text{cls}}\mathcal{L}_{\text{cls}}.$ (15)
When $\lambda_{\text{reg}}=0$, it operates as a mask matcher, whereas when $\lambda_{\text{mask}}=0$, it operates as an L1 matcher. When both $\lambda_{\text{reg}}$ and $\lambda_{\text{mask}}$ are non-zero, it is referred to as a Mask-L1 mix matcher. This hybrid matching approach, inspired by the Mask DINO framework [24], is adapted to enhance road topology understanding. Experimental results show that utilizing the instance mask formulation in both the main loss and bipartite matching (Hungarian matcher) improves performance. Furthermore, the Mask-L1 mix matcher accelerates the convergence of topology performance.
3.4 Fusion Methodology
Sensor fusion for road topology understanding has not been extensively explored in the literature. In this work, the raw features of radar and lidar point clouds are directly concatenated at the voxel space level. While similar to the BEVFusion studies [32, 41, 58] it differs by performing fusion in voxel space instead of fusing in BEV space to preserve more information. In contrast, SDMap fusion is done in the BEV space, and SDMap features are concatenated with BEV features from other sensors. The mathematical background of the proposed multi-sensor fusion pipeline is detailed in Section S.2 of the supplementary materials.
3.5 Auxiliary One-to-Many Set Prediction Loss Strategy
To improve training efficiency, an auxiliary one-to-many set prediction loss strategy [20, 35] is employed. This study, to the best of our knowledge, is the first to perform ablation experiments on this strategy for road topology understanding. Results show that this approach significantly enhances the architecture’s ability to understand road topology.
The auxiliary one-to-many set prediction loss approach involves a smaller decoder and a larger decoder with shared weights. The smaller decoder uses the ground truth set directly, while the larger decoder employs a repeated ground truth set to increase the number of positive samples. During training, both decoders are utilized; while only the smaller decoder is employed during inference. The mathematical background of this loss strategy is provided in Section S.3 of the supplementary materials.
4 Experimental Evaluation
This section provides a comprehensive analysis of the methodologies employed in this study. It commences with a detailed description of the datasets and metrics (Section 4.1). Subsequently, Section 4.2 presents an extensive experimental evaluation, encompassing an analysis of instance mask formulation, comparisons of various attention mechanisms, an examination of auxiliary one-to-many set prediction loss, an efficiency analysis of the proposed Bezier deformable attention, and a study on multi-modal fusion involving camera, radar, lidar, and SDMap. The results are juxtaposed with state-of-the-art methods on both OpenLane-V1 and OpenLane-V2.
Supplementary material provides additional details on the various loss functions (Section S.4), the implementation specifics of the proposed architecture (Section S.5), and further experiments (Section S.6) including view transformation, backbone variations, number of epochs, number of encoder and decoder layers, and efficient multi-scale implementation ablations.
4.1 Dataset and Metrics
4.1.1 Datasets
Figure 6: The PV and BEV samples from the OpenLane-V2 dataset. The top row and bottom row shows an example from Subset-A and Subset-B, respectively. (a) show the PV, highlighting centerlines and traffic elements. (b) displays the corresponding centerline instances in the BEV domain, and (c) displays centerlines with colors highlighting the topological relationships between centerlines and traffic elements. (d) demonstrate the topological relationships between different centerlines.
The OpenLane-V2 dataset [59] is a centerline detection and road topology understanding dataset and has two subsets: Subset-A and Subset-B. Subset-A is derived from the Argoverse 2 (AV2) [65] dataset, containing samples from six cities (e.g., Miami, Pittsburgh, Austin), collected using a seven-camera setup with the front camera positioned vertically. This subset contains 22,477 training samples, 4,806 validation samples, and 4,816 test samples, all with a resolution of $2048\times 1550$ pixels. The top row of Figures 6 shows the PV and BEV images of a random sample in Subset-A. The front view of the PV is cropped due to the vertical camera positioning. Multi-modal fusion studies are conducted with lidar data and SDMap and no radar data is available in Subset-A.
Subset-B is based on the NuScenes [2] dataset and contains data collected from Boston and Singapore, using a six-camera setup. It includes 27,968 training, 6,019 validation, and 6,008 test samples, with an image resolution of $1600\times 900$ pixels. Subset-B has a higher proportion of night and rainy scenes compared to Subset-A but lacks ground height information. Sensor fusion studies are implemented with both radar and lidar data but no available SDMap information in this subset. The bottom row of Figures 6 displays the PV and BEV images of a random sample from Subset-B.
The OpenLane-V1 dataset [5] is a comprehensive benchmark for 3D lane detection, derived from the Waymo Open dataset. It consists of 1000 segments with 200K frames, captured under diverse conditions at $1920\times 1280$ resolution. This dataset provides diverse and challenging scenarios for evaluating lane detection algorithms, including a wide range of weather conditions, lighting variations, and road types.
4.1.2 Metrics
For the evaluation of centerlines and traffic elements in both Subset-A and Subset-B of OpenLane-V2, the considered area extends +50 to -50 meters forward and +25 to -25 meters sideways. Both centerline detection and traffic element recognition are evaluated using the Mean Average Precision (mAP) metric. True positive samples are identified using different distance measures depending on the task. For centerline detection, the Fréchet distance and Chamfer distance are used. The Fréchet distance accounts for both distance and directionality between predicted and ground truth centerlines, whereas the Chamfer distance only considers distance, disregarding directionality. For traffic element recognition, the Intersection over Union (IoU) metric is used with a threshold of 0.75. The Fréchet distance thresholds are 1, 2, and 3 meters, while the Chamfer distance thresholds are 0.5, 1, and 1.5 meters. The mean AP of different thresholds is denoted as $\text{DET}_{l}$ for Fréchet-based mAP and $\text{DET}_{l\_ch}$ for Chamfer-based mAP, and $\text{DET}_{t}$ for IoU-based mAP.
In addition to these metrics, a specialized mAP metric is proposed to evaluate topology reasoning in the graph domain. This metric assesses the connectivity and relationships between centerlines and traffic elements. For an edge (connectivity) to be a true positive, both vertices must be correctly detected according to the Fréchet distance for centerlines and IoU criteria for traffic elements. The topology scores are defined as $\text{TOP}_{ll}$ for centerline topology and $\text{TOP}_{lt}$ for centerline-traffic element topology.
$\text{OLS}=\frac{1}{4}\bigg{[}\text{DET}_{l}+\text{DET}_{t}+f(\text{TOP}_{ll})+f(\text{TOP}_{lt})\bigg{]}.$ (16)
The evaluation framework uses the OpenLane-V2 Score (OLS) as the overall metric, which is calculated as the average of several task-specific metrics: Fréchet-based mAP for centerline prediction ($\text{DET}_{l}$), IoU-based mAP for traffic element prediction ($\text{DET}_{t}$), and the topological relationships metrics between centerlines and traffic elements ($\text{TOP}_{ll}$ and $\text{TOP}_{lt}$). The evaluation metric pipeline has been updated from version 1.0 to 1.1 to address previously identified issues, as outlined in the TopoMLP study [66].
The potential shortcomings of the 1.1 baseline are discussed in the TopoMaskV2 study [21], and a modified baseline is proposed, which is referred to as version 1.1m in this study. While the metric itself remains unchanged, the topology scores are re-mapped such that $P(x)+1\times[P(x)>0.05]$, where $P(x)$ represents the topology score. To ensure fair comparison with existing literature, state-of-the-art comparisons are conducted using V1.1, while all our ablations and hyperparameter evaluations are performed using the V1.1m baseline.
$\text{OLS}_{l}=\frac{1}{3}\bigg{[}\text{DET}_{l}+\text{DET}_{l\_ch}+f(\text{TOP}_{ll})\bigg{]}.$ (17)
To perform ablations focused exclusively on centerline prediction and centerline topology, we use the $\text{OLS}_{l}$ metric Eq. (17), which excludes components associated with traffic elements. Since the architectural framework of TopoBDA is specifically designed to improve centerline prediction, it is essential to develop metrics dedicated to centerline prediction and centerline topology prediction accuracy.
For 3D lane detection evaluation in the OpenLane-V1 dataset, the F1 metric is preferred, as it calculates the harmonic mean of precision and recall. A detection is considered a true positive if at least 75% of the compared points are within the predefined threshold of 1.5 meters or 0.5 meters from the ground truth lane dividers. The detection range for the metric is set from 3 to 103 meters in the forward direction and -10 to 10 meters in the lateral direction. A threshold of 40 meters is used to distinguish between close and far ranges, with lateral and height translation errors measured separately for these regions.
4.2 Experimental Evaluation
In this section, experimental results are presented. First, instance mask formulation is analysed, followed by a comparison of different attention mechanisms and a conceptual analysis of the auxiliary one-to-many set prediction loss. Next, the performance of sensor fusion and SDMap integration, as well as the efficiency of attention mechanisms in decoder implementations, are evaluated. Finally, the results are compared with state-of-the-art benchmarks on the OpenLane-V2 and OpenLane-V1 datasets.
The supplementary material section provides further insights into our experiments (Section S.6), including a comparative analysis of view transformation methods, a comparison of standard and efficient multi-scale implementations, an analysis of the number of encoder and decoder layers, and an evaluation of the influence of backbone, epochs, and sensor fusion on performance.
4.2.1 Analysis of Instance Mask Formulation
Table 1: Impact of indirect instance mask formulation on TopoBDA in Subset-A of OpenLane-V2 Metric with V1.1m Baseline. IMAL refers to the Instance Mask Auxiliary Loss and ML1M refers to the Mask-L1 Mix Matcher.
IMAL ML1M DETl DETl_ch TOPll OLSl
37.0 39.8 29.0 43.6
40.7 42.1 32.4 46.6
40.8 45.8 32.9 48.0
Table 1 presents the impact of indirect instance mask formulation on TopoBDA with the V1.1m metric baseline. These experiments utilize SwinB as the backbone of the architecture. The results show that incorporating the instance mask auxiliary loss (IMAL) significantly improves performance across various metrics, with a 3.7 points increase in DETl score, and a 3.3 point increase in OLSl (See Section 4.1 for the details of metrics and Eq. (17) for OLSl).
Further enhancement with the mask-L1 mix matcher (ML1M) provides additional gains, with a 3.7-point increase in DETl_ch and a 1.4-point increase in OLSl. The improvements underscore the effectiveness of the instance mask auxiliary loss and the mask-L1 mix matcher and validating their adoption for use in TopoBDA. For detailed information about the implementation of instance mask formulation please refer to Section 3.3.
4.2.2 Comparison of Different Attention Mechanisms
In Table 2(a), a comparative analysis of various attention mechanisms with the V1.1m metric baseline is presented, utilizing SwinB as the 2D backbone. The results show that deformable attention-based decoders significantly outperform both Standard Attention (SA) and Masked Attention (MA). Specifically, Single-Point Deformable Attention (SPDA) outperforms MA by 1.5 points in OLSl, but, it performs the worst among the deformable attention baselines. MPDA4 outperforms SPDA by 3.3 points in OLSl, highlighting the importance of applying the attention mechanism to different regions around the polyline instead of a single fixed center point. Notably, there is negligible performance difference between 4-point (MPDA4) and 16-point (MPDA16) multi-point deformable attentions with only a 0.1 point improvement in favor of MDPA16 in OLSl. The 4-Point Bezier Deformable Attention (BDA), which employs 4 control points, surpasses both MPDA4 and MPDA16 by 0.7 points in DETl_ch and 0.5 points in OLSl. Despite the limited performance improvement, BDA achieves these results with reduced computational complexity, as it does not require converting Bezier control points to polyline points in each transformer decoder layer using matrix multiplication, unlike MPDA. A comparison of the computational complexity of the attention mechanisms is provided in Section 4.2.5 and further details on SPDA, MPDA, and BDA are provided in Section 3.1.
Table 2: The left table presents a comparison of different attention mechanisms, Standard Attention (SA), Masked Attention (MA), Single Point Deformable Attention (SPDA), 4-point Multi-Point Deformable Attention (MPDA4), 16-point Deformable Attention (MPDA16), and Bezier Deformable Attention (BDA). The right table illustrates the impact of ground truth and query repetition (R) in the auxiliary one-to-many set prediction loss strategy, where R=0 indicates the absence of the auxiliary one-to-many set prediction loss. Both table results are on Subset-A of the Openlane-V2 dataset using the V1.1m Baseline
(a) Attention Mechanisms
Attn. DETl DETl_ch TOPll OLSl
SA 34.5 38.4 25.1 41.0
MA 35.8 40.2 26.9 42.6
SPDA 38.3 39.8 29.5 44.1
MPDA4 40.2 45.0 32.6 47.4
MPDA16 40.3 45.1 32.7 47.5
BDA 40.8 45.8 32.9 48.0
(b) Auxiliary One-to-many Set Prediction Loss
R DETl DETl_ch TOPll OLSl
0 39.4 44.0 31.4 46.5
1 39.0 45.0 32.0 46.9
2 40.1 45.0 32.1 47.2
3 40.7 45.5 32.6 47.8
4 40.8 45.8 32.9 48.0
5 41.0 45.9 33.1 48.1
4.2.3 Auxiliary One-to-Many Set Prediction Loss: Conceptual Analysis
This study evaluates the impact of varying the number of repetitions (R) of ground truths and queries on the auxiliary one-to-many set prediction loss across the key metrics: DETl, DETl_ch, TOPll, and OLSl. The results with the V1.1m metric baseline, summarized in Table 2(b), are based on experiments conducted using the SwinB backbone. The findings indicate that increasing the number of repetitions (R) enhances performance across most metrics. Consequently, with this strategy, DETl, DETl_ch, TOPll, and OLSl improve by up to 1.6, 1.9, 1.7, and 1.6 points, respectively. A detailed explanation of the auxiliary one-to-many set prediction loss strategy is provided in Section 3.5.
4.2.4 Performance Analysis of Sensor Fusion and SDMap Integration
Table 3: Sensor Fusion and SDMap Integration Ablations in Subset-A and Subset-B of OpenLane-V2 with V1.1m Baseline. The configuration with the lidar encoder (SECOND) is marked with a dagger (†) [73].
Subset Configuration DETl DETl_ch TOPll OLSl
A Camera 38.9 39.2 29.4 44.1
Camera + SDMap 42.7 48.0 35.7 50.1
Camera + Lidar 46.5 49.0 36.7 52.0
Camera + Lidar (†) 47.3 51.2 37.3 53.2
Camera + Lidar (†) + SDMap 52.0 52.8 40.0 56.0
B Camera 45.1 45.1 35.6 49.9
Camera + Radar 49.4 52.8 38.6 54.8
Camera + Lidar 57.5 59.4 46.0 61.6
Camera + Lidar (†) 57.7 60.0 46.7 62.0
Table 3 summarizes the experimental results for multi-modal fusion ablations in the OpenLane-V2 V1.1m metric baseline, with ResNet50 as the 2D backbone. Performance is evaluated across the DETl, DETl_ch, TOPll, and OLSl metrics.
The camera-only configuration serves as the baseline. In Subset-A, adding lidar significantly improves performance, with OLSl increasing from 44.1 to 52.0. Using the lidar encoder (SECOND) [73] further boosts OLSl to 53.2. Similarly, in Subset-B, integrating lidar with the camera results in substantial performance gains, with OLSl increasing from 49.9 to 61.6 and the use of the lidar encoder further increasing to 62.0.
Integration of SDMap information with the camera in Subset-A also brings notable improvements, increasing OLSl from 44.1 to 50.1. Using camera, lidar, and SDMap altogether yields the highest performance gains, achieving an OLSl of 56.0. This demonstrates the complementary nature of lidar sensors and SDMap information, providing richer contextual data and enhancing the model’s performance. Specifically, adding lidar to the camera + SDMap configuration improves OLSl from 50.1 to 56.0, while adding SDMap to the camera + lidar configuration improves OLSl from 53.2 to 56.0.
In Subset-B, integrating radar with the camera shows notable improvements, increasing OLSl from 49.9 to 54.8. However, the combination of radar and lidar does not yield further improvements beyond the camera-lidar configuration, suggesting that lidar already encapsulates the benefits provided by radar.
Figure 7 illustrates the advantages of integrating lidar and SDMap with camera sensors in the BEV domain. The corresponding camera views for these examples are shown in Figure S1 in Section S.7. In Figure 7, inaccurate centerline detections are highlighted with circular regions in the BEV images of specific modalities (camera, lidar, and SDMap) or their combinations. ‘GT’ denotes ground truth, while ‘C’, ‘SD’, and ‘L’ represent the camera, SDMap, and lidar, respectively. In the comparison figures, green centerlines denote the ground truth, while red centerlines represent the predictions, overlaid for easier comparison. Lidar proves advantageous for detecting and localizing long-distance centerlines, while SDMap improves centerline localization and orientation, and is beneficial for occluded and unseen areas. When used together, SDMap and lidar minimizes errors, offering the best best performance.
Overall, these results underscore the significant benefits of multi-modal fusion, particularly the integration of lidar and SDMap, in enhancing road topology understanding.
Figure 7: Visual demonstration in the BEV domain showing the impact of lidar and SDMap additions in Subset-A of the OpenLane-V2 dataset. C, SD, and L represent the camera, SDMap, and lidar, respectively. Green polylines indicate the ground truth and red polylines represent predictions. The circular regions highlight the inaccurate regions compared to other reference BEV images.
4.2.5 Efficiency Analysis of Attention Mechanisms in Decoder Implementations
Table 4: Decoder runtimes (in ms) for different attention types in Torch and ONNX Runtime implementations.
BDA SPDA MPDA4 MPDA16 SA MA
Torch 17.93 17.93 19.93 22.22 15.54 31.51
ONNX Runtime 9.80 9.75 9.90 10.05 14.05 38.15
The time complexity analysis of the decoder for specific attention types is presented in Table 4. For this analysis, an efficient multi-scale implementation [8], which processes different feature scales successively in each decoder layer in a round-robin fashion, is employed.
In this table, the time complexity of Bezier Deformable Attention (BDA), Single Point Deformable Attention (SPDA), Multi Point Deformable Attention (MPDA), Self Attention (SA), and Masked Attention (MA) are compared in both Torch and ONNX Runtime environments. In Torch, SA has the shortest computation duration, likely due to optimization for dynamic graphs. BDA and SPDA have similar durations, with BDA outperforming MPDA4 and MPDA16 by approximately 2ms and 4.2ms, respectively. In ONNX Runtime, SPDA has the shortest duration, followed closely by BDA with a 0.05ms difference. While the performance gap between BDA and MPDA narrows in ONNX Runtime compared to Torch, BDA consistently surpasses both MPDA4 and MPDA16. MA exhibits the highest computational complexity due to the resize operations of generated masks and foreground checks for proper attention mechanisms. Although BDA does not have the shortest computation duration, it has competitive results and it consistently outperforms other attention mechanisms in terms of main evaluation metrics (Table 2(a)).
4.2.6 Comparison with the State of the Art Results in OpenLane-V2
Table 5: Comparative Evaluation of TopoBDA Architecture with Other Methods in Subset-A and Subset-B of OpenLane-V2 with V1.1 Metric Baseline. All methods utilize the ResNet50 backbone with 24 epochs. For the sensors, C denotes the camera, L denotes the lidar, and SD denotes the SDMap.
Subset Method Sensor DETl DETt TOPll TOPlt OLS
A STSU [3] C 12.7 43.0 2.9 19.8 29.3
VectorMapNet [39] C 11.1 41.7 2.7 9.2 24.9
MapTR [34] C 8.3 43.5 2.3 8.3 24.2
TopoNet [27] C 28.6 48.6 10.9 23.9 39.8
TopoMLP [66] C 28.5 49.5 21.7 26.9 44.1
Topo2D [25] C 29.1 50.6 22.3 26.2 44.4
TopoLogic [12] C 29.9 47.2 23.9 25.4 44.1
RoadPainter [48] C 30.7 47.7 22.8 27.2 44.6
TopoFormer [47] C 34.7 48.2 24.1 29.5 46.3
TopoMaskV2 [21] C 34.5 53.8 24.5 35.6 49.4
TopoBDA (Ours) C 38.9 54.3 27.6 37.3 51.7
TopoBDA (Ours) C + L 47.3 54.0 35.5 41.9 56.4
SMERF [43] C + SD 33.4 48.6 15.4 25.4 42.9
TopoLogic [12] C + SD 34.4 48.3 28.9 28.7 47.5
RoadPainter [48] C + SD 36.9 47.1 29.6 29.5 48.2
TopoBDA (Ours) C + SD 42.7 52.4 34.3 41.7 54.6
TopoBDA (Ours) C + L + SD 52.0 52.4 38.5 45.3 58.4
B TopoNet [27] C 24.3 55.0 6.7 16.7 36.0
TopoMLP [66] C 25.2 63.1 20.7 20.3 44.7
TopoLogic [12] C 25.9 54.7 21.6 17.9 42.3
TopoFormer [47] C 34.8 58.9 23.2 23.3 47.5
TopoMaskV2 [21] C 41.6 61.1 28.7 26.1 51.8
TopoBDA (Ours) C 45.1 61.4 34.0 27.6 54.3
TopoBDA (Ours) C + L 57.7 62.9 45.0 35.2 61.7
This section presents a comparative analysis of the TopoBDA architecture against other state-of-the-art methods for road topology understanding. Table 5 illustrates this comparison for Subset-A and Subset-B of the OpenLane-V2 dataset, evaluated using the OpenLane-V2 V1.1 metric baseline.
TopoBDA achieves state-of-the-art results across all metrics in both subsets. It excels with both camera-only data and when combining camera and SDMap information, highlighting its robustness and effectiveness. Notably, TopoBDA is the best-performing model in both camera-only and camera + SDMap configurations, solidifying its position as the leading solution for road topology understanding.
The integration of lidar sensors significantly enhances performance, underscoring the importance of sensor fusion in improving road topology understanding. While incorporating SDMap information also boosts TopoBDA’s performance, leading to substantial improvements, the enhancements from lidar are more pronounced. This demonstrates the value of both lidar and map information, with lidar providing a greater impact on performance.
The highest performance gains are achieved by combining camera, lidar, and SDMap data. This synergy of multiple inputs results in top scores across all metrics, showcasing the comprehensive understanding of road topology that TopoBDA offers. Lidar sensors and SDMap provide complementary information, further enhancing the model’s performance.
4.2.7 Comparison with the State of the Art Results in OpenLane-V1
Table 6: Comparative Evaluation of TopoBDA architecture with other methods on the OpenLane-V1 Dataset. The table compares the performance comparisoın of different methods based on various metrics.
Dist. Methods Backbone F1-Score $\uparrow$ X-error X-error Z-error Z-error
near (m) $\downarrow$ far (m) $\downarrow$ near (m) $\downarrow$ far (m) $\downarrow$
1.5m PersFormer [5] ResNet-50 52.7 0.307 0.319 0.083 0.117
Anchor3DLane [19] ResNet-50 57.5 0.229 0.243 0.079 0.106
GroupLane ResNet-50 60.2 0.371 0.476 0.220 0.357
LaneCPP [52] EffNet-B7 60.3 0.264 0.310 0.077 0.117
LATR [46] ResNet-50 61.9 0.219 0.259 0.075 0.104
PVALane [80] ResNet-50 62.7 0.232 0.259 0.092 0.118
TopoBDA (Ours) ResNet-50 63.9 0.224 0.243 0.069 0.101
0.5m PersFormer [5] ResNet-50 43.2 0.229 0.245 0.078 0.106
DV-3DLane (Camera) [44] ResNet-34 52.9 0.173 0.212 0.069 0.098
LATR [46] ResNet-50 54.0 0.171 0.201 0.072 0.099
TopoBDA (Ours) ResNet-50 57.9 0.157 0.179 0.067 0.087
The proposed TopoBDA architecture outperforms other methods on the OpenLane-V1 dataset, achieving state-of-the-art results across multiple metrics (Table 6). For this analysis, we utilized the ResNet-50 backbone and processed images at 0.5×0.5$\times$0.5 × of the original resolution.
TopoBDA attains the highest F1-Score for both distance categories (1.5m and 0.5m), with scores of 63.9 and 57.9, respectively. Additionally, it records the lowest X-error and Z-error values, indicating exceptional accuracy in both lateral positioning and depth estimation.
Although LATR achieves the lowest X-error (near) for the 1.5m distance, the difference is minimal, with TopoBDA trailing by only 5 millimeters. Furthermore, TopoBDA surpasses LATR in the X-error (near) metric for the 0.5m distance, outperforming by 14 millimeters. Notably, despite the hyper-parameters being optimized for centerline detection in the OpenLane-V2 dataset, TopoBDA also achieves the best results for lane divider detection in the OpenLane-V1 dataset. Further optimization of score thresholds and hyper-parameters for this task is expected to improve TopoBDA’s performance on the OpenLane-V1 dataset.
5 Conclusion
Experimental evaluations demonstrate that TopoBDA achieves state-of-the-art results across both subsets of the OpenLane-V2 dataset, utilizing the version 1.1 metric baseline. Specifically, TopoBDA surpasses existing methodologies with a DETl score of 38.9 and an OLS score of 51.7 in Subset-A, and a DETl score of 45.1 and an OLS score of 54.3 in Subset-B. The integration of multi-modal data significantly enhances performance: fusion of camera and lidar data elevates the OLS score in Subset-A from 51.7 to 56.4, and in Subset-B from 54.3 to 61.7. The subsequent incorporation of SDMap alongside camera and lidar sensors further augments the OLS score in Subset-A from 56.4 to 58.4. These findings underscore the efficacy of TopoBDA in comprehending road topology and highlight the substantial performance improvements afforded by sensor fusion. Additionally, TopoBDA achieves superior results on the OpenLane-V1 benchmark for 3D lane detection, with F1-Scores of 63.9 for the 1.5m distance and 57.9 for the 0.5m distance.
In conclusion, this work aims to fill gaps in the literature, moving towards a comprehensive HDMap element prediction. TopoBDA represents a significant advancement in road topology understanding and 3D lane detection for autonomous driving. By leveraging Bezier deformable attention, instance mask formulation, multi-modal fusion, and an auxiliary one-to-many set prediction loss strategy, TopoBDA achieves superior performance in centerline detection and topology reasoning. This approach not only enhances computational efficiency but also sets a new benchmark in the field, highlighting its potential for practical applications in autonomous driving systems.
Acknowledgements
We acknowledge the use of the TRUBA high-performance computing infrastructure provided by TÜBİTAK ULAKBİM for the computations in this study. We also extend our gratitude to the Barcelona Supercomputing Center (BSC-CNS) and the EuroHPC Joint Undertaking for granting access to the MareNostrum 5 supercomputer, which provided additional computational resources.